you have been given the mathematical model to calculate velocity of a car accelerating from rest in straight line: v(t) = A (1-e ^ - t/tmaxspeed) (anything after the symbol ^ is to the power of) v(t is instantaneous velocity of car (m/s), t is time in seconds, tmaxspeed is time to reach maximum speed in seconds and A is constant. you are given the information that t (0-28 m/s) is 2.6s, t (400m) is 10.46s & tmaxspeed is 7s. Apply the mathematical model above to the car, use given data for the 0-28m/s and 400m times to calculate: value of the coefficient A, maximum velocity and maximum acceleration? please show graphs for this information aswell
1
Expert's answer
2020-02-27T09:35:47-0500

 "v(t) = A (1-e ^ {- t\/t_{maxspeed}})"

using the information that t (0-28 m/s) is 2.6s & tmaxspeed is 7s

"v(2.6)=28=A(1-e^{-2.6\/7}) => A=90.25m\/s"

finally "v(t)=90.25(1-e^{-t\/7})"

v(t) is ascending function, so maximum velocity will be on right end, i.e. with "t \to \infty:"

"v_{max}=v(t \to \infty)=90.25m/s"


acceleration: "a(t)=dv\/dt=90.25\/7*e^{-t\/7}"

a(t) is descending function, so the maximum acceleration will be on left end, i.e. with t=0:

"a_{max}=a(t=0)=90.25\/7=12.9m\/s^2"